OpenStudy (anonymous):

Evaluate the expression when x = –16. X-(-8)/4

jigglypuff314 (jigglypuff314):

plug in -16 for x so (-16) -(-8)/4 and simplify, but don't forget PEMDAS! :)

OpenStudy (anonymous):

(-16)-(-8)/4 (-16)-(-2) (-16) +2 -14

OpenStudy (anonymous):

–6 –2 2 6

OpenStudy (anonymous):

those are the answers

OpenStudy (anonymous):

That might change it a bit.

jigglypuff314 (jigglypuff314):

... then you might have been missing a parenthesis... try ((-16)-(-8))/4 so (-16 + 8 ) /4 (-8) /4 -2

OpenStudy (anonymous):

thanks what about this one!??

OpenStudy (anonymous):

That would have made sense if I knew the "/4" covered the whole equation.

OpenStudy (anonymous):

What is the mean of the set of numbers? {–8, –4, –3, –11} A.–6.5 B.–4.5 C.4.5 D.6.5

OpenStudy (anonymous):

Mean is green, so add and divide.

jigglypuff314 (jigglypuff314):

mean = average = (sum of numbers) / (number of numbers) (-8 + -4 + -3 + -11)/4 =

OpenStudy (anonymous):

-8+-4+-3+-11 = -26 -26/4 = -6.5

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

We're glad to help :)

OpenStudy (anonymous):

i got an 80

OpenStudy (anonymous):

Which ones did you get wrong?

OpenStudy (anonymous):

on the last one i put 6.5 instead of -6.5 haha lol

OpenStudy (anonymous):

XD Fail. IIts alright, though. Lemme know when you need help, kay?

OpenStudy (anonymous):

okay